How many hydrogen atoms are in 0.750 cm 3 of C3H8 gas at SATP?
1
Expert's answer
2020-04-23T13:06:19-0400

SATP is at 25 degrees C (298.15 K) and 101 kPa. At these conditions, the volume of 1 mol of a gas is 24.4651 liters.

n = V/Vm

V = 0.750 cm3 = 0.750 mL = 0.00075 L

n = 0.00075 / 24.4651 = 3,0656x10-5 mol

N = n * NA = 3,0656x10-5 * 6.022x1023

N = 1,846x1019 - molecules of C3H8

N = 8 * 1,846x1019 = 1,48x1020 - atoms of H